help please!
Solve the equation for x and y. x+2y=-5 x=2y-1

choices are: (5,3) (5,0) (-3,-1) (1,-3)

Answer by checkley71(8403)

X+2Y=-5
X=-2Y-5 SUBSTITUTE (-2Y-5) FOR X IN THE OTHER EQUATION
-2Y-5=2Y-1 NOW SOLVE FOR Y
-4Y=4
Y=4/-4
Y=-1 SOLUTION NOW SUBSTITUTE -1 FOR Y IN THE FIRST EQUATION
X+2Y=-5
X+2*-1=-5
X-2=-5
X=-5+2
X=-3 SOLUTION